OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to -3x + y = -1?

OpenStudy (anonymous):

Of the following...?

OpenStudy (anonymous):

Y= 1/3x + c y= -3x+c y= 1/3 + c y= 3x +c

OpenStudy (anonymous):

Thank You !

OpenStudy (anonymous):

You only have to compare the slopes, o maybe they want you to make a "graphic" of all of them and visualize the correct answer.

OpenStudy (anonymous):

Something wrong here, the slope of the given line is 3 so the slope of the perpendicular will be -1/3 but that is not one of the options (I think there is a mistake in the options)

OpenStudy (anonymous):

Oh, thats true.